# Ionization potential of radium monofluoride

### v2

@@ Abstract
The ionization potential (IP) of radium monofluoride (RaF) was measured to be 4.969(2)[10] eV, revealing a relativistic enhancement in the series of alkaline earth monofluorides. The results are in agreement with a relativistic coupled-cluster prediction of 4.969[7] eV, incorporating up to quantum electrodynamics corrections. Using the same computational methodology, an improved calculation for the dissociation energy () of 5.56[5] eV is presented. This confirms that radium monofluoride joins the small group of diatomic molecules for which , paving the way for precision control and interrogation of its Rydberg states.
@@
Molecules containing isotopes of radium have been proposed as being promising systems in which to study the fundamental symmetries of the Universe [8, 9], particularly in the hadronic sector of the Standard Model, where the rare octupole deformation of certain radium isotopes significantly boosts their sensitivity to -violating phenomena [10].
A notable example is the RaF molecule, which was theoretically predicted [8] and later experimentally confirmed to be directly laser-coolable [11, 12, 13]. Leveraging the unique properties of high-lying Rydberg states of RaF through manipulating them with external fields, could enable the sensitive control of these molecules [14, 15], offering complementary opportunities for future precision measurements. The realization of such techniques however relies upon the condition that .
@@
The resulting thresholds are shown in Fig. 1 b) and c) for the two-step (teal squares) and three-step schemes (blue circles), respectively. Using the same method as presented in Ref. [19], the two sets of data were fit using a Sigmoid function and the IPs determined from the energy at which the ion rate saturates. This was done by determining the energy at which the tangent to the inflection point of the fitted Sigmoid curves intersects the maximum ion rate, and its corresponding boundaries. The IPs and their statistical uncertainties are shown as the vertical dashed-and-dotted lines and the narrow dark grey bands in Fig. 1 b) and c). A systematic uncertainty from the fitting procedure is assigned by taking the energy difference between the inflection points of the curves and the extracted IPs. This is due to the lack of consistency in the literature regarding where the IP lies with respect to measured ionization thresholds. The total combined statistical and systematic uncertainties are shown as the wide light gray bands in Fig. 1 b) and c). More details on the fitting and its justification can be found in the Supplementary Material. The ionization threshold measured using the three-step scheme (Fig. 1 c)) can be seen to increase more sharply when compared to the two-step scheme (Fig. 1 b)). This is due to the additional rotational angular momentum () selectivity introduced by having two sequential resonant excitations before the ionization step in the three-step scheme. This in turn reduced the number of high-lying Rydberg states populated by the ionization laser at energies below the IP, resulting in a sharper threshold. The ionization thresholds were determined to be 4.966(3)[10] eV and 4.970(1)[2] eV for the two-step and three-step scheme, respectively, with the statistical errors shown in curved brackets and the systematic errors shown in straight brackets.
More details on the fitting and its justification can be found in the Supplementary Material.
@@
CBS-DC-CCSD111ZPE correction is included. | 4.932 | 5.454 |
CBS-DC-CCSD(T)111ZPE correction is included. | 4.983 | 5.547 |
+aug+ae.vs.cv | 4.977 | 5.554 |
+T | 4.978 | 5.549 |
+Breit | 4.976 | 5.547 |
+QED | 4.969 | 5.556 |
Theoretical | 4.969[7] | 5.56[5] |
Experimental | 4.969(2)[10] | 5.64(24)222Scaled from Ref. [12]. |
@@
The various higher-order corrections are added to the adiabatic CBS-DC-CCSD(T) IP and the final theoretical value was determined to be 4.969[7] eV as shown in Table 2. Using the same computational method, the dissociation energy of RaF was calculated to be 5.56[5] eV. The uncertainty on the calculated IP and was evaluated through further computational investigation using similar procedures that were previously employed for various properties of both atoms and molecules [31, 32, 33, 34]. The details of the uncertainty treatment for both properties can be found in the Supplemental Material.
@@
The previously calculated values of the dissociation energy, , for RaF vary between 3.980 eV [8] and 5.282 eV [46] with no corresponding uncertainties reported. Using the same computational method as for the IP, which achieves a more extensive and rigorous treatment of higher-order contributions compared to previous studies, an improved calculation of the RaF equal to 5.56[5] eV is presented. This is in agreement with the RaF extracted from Morse potential fits in Ref. [12], provided they are scaled by 1.55(6). Dissociation energy determinations from Morse potential fits use linear extrapolations that significantly underestimate the true dissociation energy in molecules where ionic bonding is important [47, 48]. By taking the extrapolated values for CaF [49], SrF [50] and BaF [51], and comparing them to the true measured dissociation energies [52], the aforementioned scaling factor is obtained, which results in a RaF of 5.64(24) eV (using the value from Ref. [12]). The resulting value and uncertainty of the computed , corroborated by the agreement between experiment and theory for the IP, confirms that RaF joins BaF in the rare class of diatomic molecules for which IP.
Conclusion: The ionization potential of radium monofluoride (RaF) was measured with the CRIS experiment at CERN-ISOLDE to be 4.969(2)[10] eV, through consistent observations of the ionization thresholds starting from two different excited electronic states. The experimental results were compared against predictions calculated with single-reference relativistic coupled-cluster theory that incorporates triple excitations, basis-set, Breit-effect and QED corrections. A good agreement is obtained between experiment and theory for RaF, in addition to its lighter homologues. The dissociation energy was found to lie above the IP, offering the possibility of accessing and utilizing Rydberg states in RaF, without loss due to pre-dissociation. This opens up several opportunities for controlling and manipulating the molecule via external fields [14, 15]. Additionally, these states can also offer an alternative window into probing the structure of radium nuclei [5].

### v3

@@ Abstract
The ionization potential (IP) of radium monofluoride (RaF) was measured to be 4.969(2)[10] eV, revealing a relativistic enhancement in the series of alkaline earth monofluorides. The results are in agreement with a relativistic coupled-cluster prediction of 4.969[7] eV, incorporating up to quantum electrodynamics corrections. Using the same computational methodology, an improved calculation for the dissociation energy () of 5.54[5] eV is presented. This confirms that radium monofluoride joins the small group of diatomic molecules for which , paving the way for precision control and interrogation of its Rydberg states.
@@
Molecules containing isotopes of radium have been proposed as being promising systems in which to study the fundamental symmetries of the Universe [8, 9, Wilkins2023], particularly in the hadronic sector of the Standard Model, where the rare octupole deformation of certain radium isotopes significantly boosts their sensitivity to -violating phenomena [10]. A notable example is the RaF molecule, which was theoretically predicted [8] and later experimentally confirmed to be directly laser-coolable [11, 12, 13]. Leveraging the unique properties of high-lying Rydberg states of RaF through manipulating them with external fields, could enable the sensitive control of these molecules [14, 15], offering complementary opportunities for future precision measurements. The realization of such techniques however relies upon the condition that .
@@
The resulting thresholds are shown in Fig. 1 b) and c) for the two-step (teal squares) and three-step schemes (blue circles), respectively. Using the same method as presented in Ref. [19], the two sets of data were fit using a Sigmoid function and the IPs determined from the energy at which the ion rate saturates. This was done by determining the energy at which the tangent to the inflection point of the fitted Sigmoid curves intersects the maximum ion rate, and its corresponding boundaries. The IPs and their statistical uncertainties are shown as the vertical dashed-and-dotted lines and the narrow dark grey bands in Fig. 1 b) and c). A systematic uncertainty from the fitting procedure is assigned by taking the energy difference between the inflection points of the curves and the extracted IPs. This is due to the lack of consistency in the literature regarding where the IP lies with respect to measured ionization thresholds. The total combined statistical and systematic uncertainties are shown as the wide light gray bands in Fig. 1 b) and c). More details on the fitting and its justification can be found in the Supplementary Material. The ionization threshold measured using the three-step scheme (Fig. 1 c)) can be seen to increase more sharply when compared to the two-step scheme (Fig. 1 b)). This is due to the additional rotational angular momentum () selectivity introduced by having two sequential resonant excitations before the ionization step in the three-step scheme. This in turn reduced the number of high-lying Rydberg states populated by the ionization laser at energies below the IP, resulting in a sharper threshold. The ionization thresholds were determined to be 4.966(3)[10] eV and 4.970(1)[2] eV for the two-step and three-step scheme, respectively, with the statistical errors shown in curved brackets and the systematic errors shown in straight brackets. More details on the fitting and its justification can be found in the Supplementary Material.
@@
CBS-DC-CCSD | 4.932111ZPE correction is included. | 5.427 |
CBS-DC-CCSD(T) | 4.983111ZPE correction is included. | 5.520 |
+aug+ae.vs.cv | 4.977 | 5.538 |
+T | 4.978 | 5.534 |
+Breit | 4.976 | 5.532 |
+QED | 4.969 | 5.541 |
Theoretical | 4.969[7] | 5.54[5] |
Experimental | 4.969(2)[10] | 5.61(24)222Scaled from Ref. [12]. |
@@
The various higher-order corrections are added to the adiabatic CBS-DC-CCSD(T) IP and the final theoretical value was determined to be 4.969[7] eV as shown in Table 2. Using the same computational method, the dissociation energy of RaF was calculated to be 5.54[5] eV. The uncertainty on the calculated IP and was evaluated through further computational investigation using similar procedures that were previously employed for various properties of both atoms and molecules [31, 32, 33, 34]. The details of the uncertainty treatment for both properties can be found in the Supplemental Material.
@@
The previously calculated values of the ZPE-corrected dissociation energy, , for RaF vary between 3.954 eV [8] and 5.255 eV [17] with no corresponding uncertainties reported. Using the same computational method as for the IP, which achieves a more extensive and rigorous treatment of higher-order contributions compared to previous studies, an improved calculation of the RaF equal to 5.54[5] eV is presented. This is in agreement with the RaF extracted from Morse potential fits in Ref. [12], provided they are scaled by 1.55(6). Dissociation energy determinations from Morse potential fits use linear extrapolations that significantly underestimate the true dissociation energy in molecules where ionic bonding is important [46, 47]. By taking the extrapolated values for CaF [48], SrF [49] and BaF [50], and comparing them to the true measured dissociation energies [51], the aforementioned scaling factor is obtained, which results in a RaF of 5.61(24) eV (using the value from Ref. [12]). The resulting value and uncertainty of the computed , corroborated by the agreement between experiment and theory for the IP, confirms that RaF joins BaF in the rare class of diatomic molecules for which IP.
Conclusion: The ionization potential of radium monofluoride (RaF) was measured with the CRIS experiment at CERN-ISOLDE to be 4.969(2)[10] eV, through consistent observations of the ionization thresholds starting from two different excited electronic states. The dissociation energy was found to lie above the IP, offering the possibility of accessing and utilizing Rydberg states in RaF, without loss due to pre-dissociation. This opens up several opportunities for controlling and manipulating the molecule via external fields [14, 15]. Additionally, these states can also offer an alternative window into probing the structure of radium nuclei [5].